# Enhancing Vision-Language Pre-training with Rich Supervisions

### v1

@@ 1 Introduction
In recent years, there has been significant progress in Language Models (LMs) [7, 48, 11, 43] and Vision Language Models (VLMs) [56, 46, 10, 40, 37, 69, 24, 45, 73, 35, 2, 72, 6, 58, 19, 74, 70, 57, 23, 41, 65, 22, 61, 68, 67, 63, 60, 75, 16, 62, 52, 8, 15, 1, 26, 25, 55, 36, 66, 64] exhibiting strong zero-shot generalization and adaptability to a wide range of tasks. Though they may differ in architecture, data and task formulation, such foundational models predominantly rely on large-scale pre-training on giant corpora of web scraped data which serves as the source of generalization capability - C4 [47], The Pile [18], Laion 5B [50].
@@ 3.2.2 Optical Character Recognition - OCR
The OCR objective aims to train the model with the ability for spatial understanding. It takes in a screenshot with drawn bounding box specifying a region, and outputs a “word<x><y><x><y>word<x><y><x><y>…” sequence. To limit the sequence length, we choose bounding box for a random region that contains at most 50 words. The OCR objective empowers the model with the ability to spatially understand the screenshot, bringing benefits for general detection and grounding tasks.
@@ 3.2.7 Table Detection
To closely mimic the downstream task for table detection, we construct table detection on our screenshot data. The construction is as simple as merging the bounding box for the elements with <table[id]> contained in their Xpaths, which results in the ground truth bounding box for each table. We then ask the model to predict the following sequence:<x><y><x><y><x> <y><x><y>...
@@ 3.3 Architecture
| Methods | Pre-training | Pre-training | Finetune | ChartQA | RefExp | Widget | Screen | WebSRC |
@@
| S4* (Ours) | S4 Data - 2M | S4 | 8 | 50.5 | 92.4 | 130.5 | 103.2 | 60.5 |
| S4* (Ours) | S4 Data - 15M | S4 | 8 | 55.0 | 94.9 | 130.6 | 105.7 | 61.1 |
@@ 4.1 Implementation Details
Pre-training schema. We propose 2 pre-training schemes, S4 for natural language generation and S4 for localization, targeting on different downstream tasks. Specifically, S4 includes the baseline screen parsing task and all the tasks on natural language generation, including Attribute Prediction, Table Parsing, Title Generation, and Node Relation Prediction. S4 comprises of tasks with bounding box generations, including OCR, Image Grounding, Element Grounding, Table Detection and Layout Analysis, in addition to the screen parsing task. During pre-training, we randomly sample one task for each image with uniform distribution.
@@ 4.1.1 Pretraining Settings
We conducted pretraining on both 2 million and 15 million subsets of our S4 dataset, during which we set the screenshot’s viewport to 1280*1280. We initialize our model with weights from Pix2struct-base and set our batch size to 32 for each node and use 4 A100 nodes during pretraining. The maximum sequence length for the pretraining targets is 128 and the patch size for the input image is 2048. Our optimizer is AdamW with learning rate set to 1e-4 with cosine decay, and for both 2M and 15M subsets we pretrain with 1 epoch per pretraining task. For instance, for S4 pretraining with the 2M subset, there are 5 tasks so the total training sample is 5*2M = 10M. Note that since for each screenshots we can obtain multiple tasks, the models sees the same subset of screenshots regardless of the number of pretraining tasks.
@@ 4.2 Chart, Web, and UI Understanding
| (Table Det.) | (Modern Subset) | |||||
@@
| S4* (Ours) | S4 Data - 2M | S4 | 83.6 | 92.5 | 98.4 | 70.7 |
| S4* (Ours) | S4 Data - 15M | S4 | 84.3 | 93.1 | 99.0 | 79.4 |
@@ 4.2.1 Datasets
UI RefExp UI Referential Expression (RefExp) [4] is a dataset specifically designed for grounding referring expressions for UI elements in screenshots. Current SOTA usually approach this problem in a simplified classification formulation: given a question and a candidate widget from annotation, the model is asked to predict whether the widget and question are related. This setting requires little localization ability from the model as the candidates widget bounding boxes are provided as inputs. We call this classification setting RefExpand report the classification accuracy as the metric.
@@ 4.2.2 Settings
Following the same training and evaluation protocol [32], our model was pre-trained with S4 objectives and finetuned on the Chart, Web, and UI Understanding tasks. Since there’s no access to Google’s 80M private data, we compare to two PixStruct variations. The first one is with the weights pre-trained on its private data using screen parsing released by the original author. The second one is initialized with the former’s weights, and is further pre-trained on our 2M and 15M S4 data with only screen parsing objective. To have a fair comparison, our model is initialized with the same weights, and pre-trained on the same amount of data (2M and 15M S4 data) but with extra tasks. We also compare to Donut[27], which is another model uses pure-vision inputs and produces text predictions.
@@ 4.2.3 Results
We tabulate the results in Tab. 1. Our method consistently outperforms Pix2Struct on all downstream tasks with significant margins when pre-trained with the same data. Specifically, when pre-trained with 15M image-text pairs, our method achieves 2.9, 6.8, 1.4, 1.2, and 1.0 improvement over Pix2Struct on ChartQA, RefExp, Widget Captioning, Screen Summarization, and WebSRC, respectively. Notice that the largest improvement is obtained on RefExp, because the pre-training tasks we proposed, such the attribute prediction and node relation prediction, help the model build robust connections between the UI elements and their referring expressions. In addition, when more data (15M) is available, our pre-training scheme S4 gets improved accuracy compared to less training data (2M). The outstanding results comparing to the baseline and Donut demonstrate the efficacy of the proposed S4 pre-training scheme for downstream tasks that involve chart, web, and UI understanding.
@@ 4.3 Detection and Grounding
| Titling | Attibute Pred. | NRP | Table Parsing. | Screen Parsing | ChartQA | Widget Cap. | Screen Sum. | RefExp |
@@
| Table Detection | Layout Analysis | Image & Element Grounding | OCR | Screen Parsing | ICDAR | RefExp |
@@ 4.3.1 Datasets
UI RefExp As mentioned earlier, current works mostly treat the UI RefExp task as a binary classification problem using the ground truth bounding boxes as candidates, making it less challenging as it does not measure whether the model can localize the UI elements. In this work, we propose a new task, UI RefExp, to include the element grounding into the challenges. In this task, the input is only the screenshot with the text description, and the model is asked to predict the bounding box of the related UI element directly, thus ”candidate free”. For evaluation, the predicted bounding box will be matched to the closest ground truth box to compute the accuracy.
@@ 4.3.2 Settings
Our model was pre-trained with S4 for the benefits on localization related tasks. The model is then fine-tuned and evaluated on each downstream task dataset. We compare to two PixStruct variations. The first one is the weights pre-trained on its private data using screen parsing released by the original author. The second one is initialized with the former weights and is further pre-trained on our S4 data 2M with only screen parsing, to be compared to our model pre-trained on the same amount of data.
@@ 4.3.3 Results
The evaluation results on all detection and grounding related tasks are tabulated in Tab. 2. Our method shows clear advantages over the baseline Pix2Struct model that was only pre-trained with the screen parsing task. Specifically, on RefExp, when pre-trained with 2M data, our method outperforms Pix2Struct by a significant margin of 30.9 (83.6 vs. 52.7). This is because our pre-training tasks like OCR prediction and element grounding, help the model learn to localize items in the image given their semantic descriptions. Similarly, on ICDAR, which only has 600 training samples, our method achieves 70.7 AP with 2M pre-training data, while the baseline Pix2Struct only obtains 3.3, due to its lack of localization ability. When there are more fine-tuning data for the downstream tasks (PublayNet 1M & PubTables 400K), Pix2Struct can learn to localize the objects and obtain decent performance, but our training scheme S4 still benefits the model and improves the performance by 1.5 on PublayNet and 1.3 on PubTables.
The benefits of S4 pre-training becomes more prominent when more pre-train data is available. Our model pre-trained with 15M data consistently improves the accuracy on all four downstream tasks compared to 2M pre-train data. In particular, on ICDAR the 15M pre-train data improves the accuracy from 70.7 to 79.4, showing that the pre-training task benefits more when the downstream task has less data. It is worth noting that, as a generic auto-regressive text generation model, our method with 15M pre-training data achieves comparable performance on PublayNet and PubTables to detection specific models like DeTR and Dit-B, showing that sufficient pre-training with proper tasks helps close the gap between auto-regressive models and detection-specific architectures.
@@ 4.4 Contribution of each task
We conducted ablative studies to show the effectiveness of each individual pre-training tasks besides screen parsing. For S4, we evaluate on ChartQA, Widget Captioning, Screen Summarization, and RefExp, by adding the natural language related tasks gradually. For S4, we also add the localization related tasks incrementally and evaluate on RefExp and ICDAR. The results are shown in Tab. 3. Observe that the downstream task usually benefits from the addition of the most related pre-training task. For example, Screen Summarization gets 2.1 performance improvement when the screen tilting pre-training task is added, while the other tasks have little effect on the performance. The attribute prediction task encourages the model to associate website elements to their text description. Therefore, adding it to the pre-training scheme significantly improves the performance on both Widget Captioning and RefExp, which requires the model to associate UI elements to texts. Similarly, adding all the localization related pre-train task substantially improves the model’s ability on grounding elements, resulting in higher accuracy on both ICDAR and RefExp.
@@ Appendix A Quantitative Results on Joint Training
We provide some additional discussions for the impact of joint training language and location tasks. Directly combining S4 objectives and S4 objectives, which we denote as S4, harms the performance on most downstream tasks. For S4, we use the following weighted loss as it gives the best average performance across all tasks:
@@
| Methods | Pre-training | Pre-training | Finetune | ChartQA | RefExp | Widget | Screen | WebSRC |
@@
| S4* (Ours) | S4 Data - 15M | S4 | 8 | 55.0 | 94.9 | 130.6 | 105.7 | 61.1 |
| S4* (Ours) | S4 Data - 15M | S4 | 8 | 52.1 | 94.9 | 128.4 | 101.5 | 60.3 |
@@
| (Table Det.) | (Modern Subset) | |||||
@@
| S4* (Ours) | S4 Data - 15M | S4 | 84.3 | 93.1 | 99.0 | 79.4 |
| S4* (Ours) | S4 Data - 15M | S4 | 79.2 | 91.6 | 97.7 | 76.7 |

### v2

@@ 1 Introduction
In recent years, there has been significant progress in Language Models (LMs) [7, 48, 11, 43] and Vision Language Models (VLMs) [57, 46, 10, 52, 40, 37, 70, 24, 45, 74, 35, 2, 73, 6, 59, 19, 75, 71, 58, 23, 41, 66, 22, 62, 69, 68, 64, 61, 76, 16, 63, 53, 8, 15, 1, 26, 25, 56, 36, 67, 65] exhibiting strong zero-shot generalization and adaptability to a wide range of tasks. Though they may differ in architecture, data and task formulation, such foundational models predominantly rely on large-scale pre-training on giant corpora of web scraped data which serves as the source of generalization capability - C4 [47], The Pile [18], Laion 5B [50].
@@ 3.2.2 Optical Character Recognition - OCR
The OCR objective aims to train the model with the ability for spatial understanding. It takes in a screenshot with drawn bounding box specifying a region, and outputs a “word0<x0><y0><x0><y0>word1<x1><y1><x1><y1>…” sequence. To limit the sequence length, we choose bounding box for a random region that contains at most 50 words. The OCR objective empowers the model with the ability to spatially understand the screenshot, bringing benefits for general detection and grounding tasks.
@@ 3.2.7 Table Detection
To closely mimic the downstream task for table detection, we construct table detection on our screenshot data. The construction is as simple as merging the bounding box for the elements with <table[id]> contained in their Xpaths, which results in the ground truth bounding box for each table. We then ask the model to predict the following sequence:<xtable0><ytable0><xtable0><ytable0><xtable1> <ytable1><xtable1><ytable1>...
@@ 3.3 Architecture
| Methods | Pre-training | Pre-training | Finetune | ChartQA | RefExpcls | Widget | Screen | WebSRC |
@@
| S4* (Ours) | S4 Data - 2M | S4NL | 8 | 50.5 | 92.4 | 130.5 | 103.2 | 60.5 |
| S4* (Ours) | S4 Data - 15M | S4NL | 8 | 55.0 | 94.9 | 130.6 | 105.7 | 61.1 |
@@ 4.1 Implementation Details
Pre-training schema. We propose 2 pre-training schemes, S4NL for natural language generation and S4Loc for localization, targeting on different downstream tasks. Specifically, S4NL includes the baseline screen parsing task and all the tasks on natural language generation, including Attribute Prediction, Table Parsing, Title Generation, and Node Relation Prediction. S4Loc comprises of tasks with bounding box generations, including OCR, Image Grounding, Element Grounding, Table Detection and Layout Analysis, in addition to the screen parsing task. During pre-training, we randomly sample one task for each image with uniform distribution.
@@ 4.1.1 Pretraining Settings
We conducted pretraining on both 2 million and 15 million subsets of our S4 dataset, during which we set the screenshot’s viewport to 1280*1280. We initialize our model with weights from Pix2struct-base and set our batch size to 32 for each node and use 4 A100 nodes during pretraining. The maximum sequence length for the pretraining targets is 128 and the patch size for the input image is 2048. Our optimizer is AdamW with learning rate set to 1e-4 with cosine decay, and for both 2M and 15M subsets we pretrain with 1 epoch per pretraining task. For instance, for S4NL pretraining with the 2M subset, there are 5 tasks so the total training sample is 5*2M = 10M. Note that since for each screenshots we can obtain multiple tasks, the models sees the same subset of screenshots regardless of the number of pretraining tasks.
@@ 4.2 Chart, Web, and UI Understanding
| cand_free | (Table Det.) | (Modern Subset) | ||||
@@
| S4* (Ours) | S4 Data - 2M | S4Loc | 83.6 | 92.5 | 98.4 | 70.7 |
| S4* (Ours) | S4 Data - 15M | S4Loc | 84.3 | 93.1 | 99.0 | 79.4 |
@@ 4.2.1 Datasets
UI RefExpcls UI Referential Expression (RefExp) [4] is a dataset specifically designed for grounding referring expressions for UI elements in screenshots. Current SOTA usually approach this problem in a simplified classification formulation: given a question and a candidate widget from annotation, the model is asked to predict whether the widget and question are related. This setting requires little localization ability from the model as the candidates widget bounding boxes are provided as inputs. We call this classification setting RefExpclsand report the classification accuracy as the metric.
@@ 4.2.2 Settings
Following the same training and evaluation protocol [32], our model was pre-trained with S4NL objectives and finetuned on the Chart, Web, and UI Understanding tasks. Since there’s no access to Google’s 80M private data, we compare to two PixStruct variations. The first one is with the weights pre-trained on its private data using screen parsing released by the original author. The second one is initialized with the former’s weights, and is further pre-trained on our 2M and 15M S4 data with only screen parsing objective. To have a fair comparison, our model is initialized with the same weights, and pre-trained on the same amount of data (2M and 15M S4 data) but with extra tasks. We also compare to Donut[27], which is another model uses pure-vision inputs and produces text predictions.
@@ 4.2.3 Results
We tabulate the results in Tab. 1. Our method consistently outperforms Pix2Struct on all downstream tasks with significant margins when pre-trained with the same data. Specifically, when pre-trained with 15M image-text pairs, our method achieves 2.9, 6.8, 1.4, 1.2, and 1.0 improvement over Pix2Struct on ChartQA, RefExpcls, Widget Captioning, Screen Summarization, and WebSRC, respectively. Notice that the largest improvement is obtained on RefExpcls, because the pre-training tasks we proposed, such the attribute prediction and node relation prediction, help the model build robust connections between the UI elements and their referring expressions. In addition, when more data (15M) is available, our pre-training scheme S4NL gets improved accuracy compared to less training data (2M). The outstanding results comparing to the baseline and Donut demonstrate the efficacy of the proposed S4 pre-training scheme for downstream tasks that involve chart, web, and UI understanding.
@@ 4.3 Detection and Grounding
| Titling | Attibute Pred. | NRP | Table Parsing. | Screen Parsing | ChartQA | Widget Cap. | Screen Sum. | RefExpcls |
@@
| Table Detection | Layout Analysis | Image & Element Grounding | OCR | Screen Parsing | ICDAR | RefExpcand_free |
@@ 4.3.1 Datasets
UI RefExpcand_free As mentioned earlier, current works mostly treat the UI RefExp task as a binary classification problem using the ground truth bounding boxes as candidates, making it less challenging as it does not measure whether the model can localize the UI elements. In this work, we propose a new task, UI RefExpcand_free, to include the element grounding into the challenges. In this task, the input is only the screenshot with the text description, and the model is asked to predict the bounding box of the related UI element directly, thus "candidate free". For evaluation, the predicted bounding box will be matched to the closest ground truth box to compute the accuracy.
@@ 4.3.2 Settings
Our model was pre-trained with S4Loc for the benefits on localization related tasks. The model is then fine-tuned and evaluated on each downstream task dataset. We compare to two PixStruct variations. The first one is the weights pre-trained on its private data using screen parsing released by the original author. The second one is initialized with the former weights and is further pre-trained on our S4 data 2M with only screen parsing, to be compared to our model pre-trained on the same amount of data.
@@ 4.3.3 Results
The evaluation results on all detection and grounding related tasks are tabulated in Tab. 2. Our method shows clear advantages over the baseline Pix2Struct model that was only pre-trained with the screen parsing task. Specifically, on RefExpcand_free, when pre-trained with 2M data, our method outperforms Pix2Struct by a significant margin of 30.9 (83.6 vs. 52.7). This is because our pre-training tasks like OCR prediction and element grounding, help the model learn to localize items in the image given their semantic descriptions. Similarly, on ICDAR, which only has 600 training samples, our method achieves 70.7 AP with 2M pre-training data, while the baseline Pix2Struct only obtains 3.3, due to its lack of localization ability. When there are more fine-tuning data for the downstream tasks (PublayNet 1M & PubTables 400K), Pix2Struct can learn to localize the objects and obtain decent performance, but our training scheme S4Loc still benefits the model and improves the performance by 1.5 on PublayNet and 1.3 on PubTables.
The benefits of S4Loc pre-training becomes more prominent when more pre-train data is available. Our model pre-trained with 15M data consistently improves the accuracy on all four downstream tasks compared to 2M pre-train data. In particular, on ICDAR the 15M pre-train data improves the accuracy from 70.7 to 79.4, showing that the pre-training task benefits more when the downstream task has less data. It is worth noting that, as a generic auto-regressive text generation model, our method with 15M pre-training data achieves comparable performance on PublayNet and PubTables to detection specific models like DeTR and Dit-B, showing that sufficient pre-training with proper tasks helps close the gap between auto-regressive models and detection-specific architectures.
@@ 4.4 Contribution of each task
We conducted ablative studies to show the effectiveness of each individual pre-training tasks besides screen parsing. For S4NL, we evaluate on ChartQA, Widget Captioning, Screen Summarization, and RefExpcls, by adding the natural language related tasks gradually. For S4Loc, we also add the localization related tasks incrementally and evaluate on RefExpcand_free and ICDAR. The results are shown in Tab. 3. Observe that the downstream task usually benefits from the addition of the most related pre-training task. For example, Screen Summarization gets 2.1 performance improvement when the screen tilting pre-training task is added, while the other tasks have little effect on the performance. The attribute prediction task encourages the model to associate website elements to their text description. Therefore, adding it to the pre-training scheme significantly improves the performance on both Widget Captioning and RefExpcls, which requires the model to associate UI elements to texts. Similarly, adding all the localization related pre-train task substantially improves the model’s ability on grounding elements, resulting in higher accuracy on both ICDAR and RefExpcand_free.
@@ Appendix A Quantitative Results on Joint Training
We provide some additional discussions for the impact of joint training language and location tasks. Directly combining S4NL objectives and S4Loc objectives, which we denote as S4Joint, harms the performance on most downstream tasks. For S4Joint, we use the following weighted loss as it gives the best average performance across all tasks:
@@
| Methods | Pre-training | Pre-training | Finetune | ChartQA | RefExpcls | Widget | Screen | WebSRC |
@@
| S4* (Ours) | S4 Data - 15M | S4NL | 8 | 55.0 | 94.9 | 130.6 | 105.7 | 61.1 |
| S4* (Ours) | S4 Data - 15M | S4Joint | 8 | 52.1 | 94.9 | 128.4 | 101.5 | 60.3 |
@@
| cand_free | (Table Det.) | (Modern Subset) | ||||
@@
| S4* (Ours) | S4 Data - 15M | S4Loc | 84.3 | 93.1 | 99.0 | 79.4 |
| S4* (Ours) | S4 Data - 15M | S4Joint | 79.2 | 91.6 | 97.7 | 76.7 |